OpenStudy (anonymous):

what is 2x^2-9x+9=0 using the quadratic formula?

OpenStudy (apoorvk):

Hmm compare your quadratic equation to the standard equation \(ax^2 + bx + c=0\), and find out 'a', 'b' and 'c' for your particular equation. Now the quadratic formula is: \[\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Substitute obtained values of 'a', 'b' and 'c', calculate and enjoy!

OpenStudy (anonymous):

i dont know if its right or wrong \

OpenStudy (apoorvk):

what's R/W? Do you seem to be getting an answer?